OpenStudy (anonymous):

From a class of 14 boys and 9 girls, how many ways can I choose a committee of 6 to analyze classroom productivity with and equal number of boys and girls? a) 30,576 b)252,252 c)48 d)126

OpenStudy (anonymous):

Don't quote me on this but I think it is D. 126....

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Your Welcome, good Luck! :)

OpenStudy (anonymous):

for some reason i think its choice a but I'm not 100% sure.can anyone else help?

OpenStudy (anonymous):

ASAP

OpenStudy (anonymous):

please anyone help asap. I don't need a too detailed answer

OpenStudy (kropot72):

The number of ways to choose 3 girls from 14 girls is 14C3. Each of these ways can be combined with each of the 9C3 ways of choosing 3 boys from 9 boys. Therefore the total number of ways of choosing 6 students with equal numbers of boys and girls is given by: \[\large 14C3\times9C3=\frac{14\times13\times12}{6}\times\frac{9\times8\times7}{6}=you\ can\ calculate\]

OpenStudy (anonymous):

that's not one of the answer choices which is odd because it looks right

OpenStudy (kropot72):

What did you get for each of the fractions?

OpenStudy (anonymous):

i got the same as you to be honest but it wasn't one of the answer choices so i thought i did it wrong

OpenStudy (kropot72):

But what did you get for each of the fractions?